### v1

# The Color-Clinical Decoupling: Why Perceptual Calibration Fails Clinical Biomarkers in Smartphone Dermatology

Abstract
Background: The rapid expansion of smartphone-based tele-dermatology assumes that colorimetric calibration—minimizing perceptual color error ($\Delta E$)—ensures clinical reliability. This assumption remains untested at scale, particularly for underrepresented skin phototypes (Fitzpatrick III–IV).
Objective: To investigate whether standard colorimetric calibration translates to clinical biomarker reliability, and to identify the dominant sources of variance in cross-device skin color measurement.
Methods: We analyzed 43,425 images from 965 Korean subjects (Fitzpatrick III–IV) across DSLR, tablet, and smartphone devices using the AI Hub Korean Skin Condition Measurement dataset. Linear Color Correction Matrix (CCM) normalization—the industry standard—was applied to align consumer devices to a DSLR reference. Performance was assessed using CIEDE2000 color differences ($\Delta E$) and Intraclass Correlation Coefficients (ICC) for Melanin Index, Erythema Index, and Individual Typology Angle (ITA).
[…]
Conclusions: Standard colorimetric calibration creates false confidence: perceptually accurate images can yield unreliable clinical biomarkers. The ITA formula’s disproportionate sensitivity to $b^{*}$ channel errors—the noisiest color component—explains this decoupling mathematically. The dominance of anatomical variance over device variance challenges single-patch calibration assumptions and demands region-aware protocols. These findings establish that current smartphone dermatology standards are insufficient for clinical-grade biomarker extraction.
Keywords: tele-dermatology, color-clinical decoupling, smartphone validation, ITA, skin colorimetry, calibration failure, Fitzpatrick skin type
1 Introduction
The proliferation of smartphone-based dermatological applications has created unprecedented opportunities for remote skin health monitoring.[Esteva2017, Haenssle2018] Deep learning systems have demonstrated expert-level performance in melanoma classification,[Marchetti2018, Brinker2019] fueling expectations that consumer devices could enable accessible platforms for quantitative skin analysis. Central to this vision is an implicit assumption: that colorimetric calibration—the minimization of perceptual color error ($\Delta E$)—ensures clinical reliability for derived biomarkers.
This assumption has never been rigorously tested.
The optical complexity of human skin presents fundamental challenges for camera-based colorimetry. Skin functions as a multilayered turbid medium, with color determined by the absorption spectra of melanin in the epidermis and hemoglobin in the dermal vasculature.[Jacques2013, Anderson1981] When this continuous spectral information is sampled by RGB sensors, the fine spectral detail required to separate chromophore contributions is irreversibly lost.[Zonios2001, Nishidate2004] Device-specific variations compound this limitation: different sensors produce inconsistent RGB values for identical skin samples due to variations in spectral sensitivities and image signal processing (ISP) pipelines.[Finlayson2015, Karaimer2016]
The standard solution—linear Color Correction Matrix (CCM) normalization—transforms device RGB values toward a reference standard by minimizing the Euclidean distance in color space.[Hong2001, Cheung2004] This approach optimizes for perceptual color fidelity, as measured by the CIEDE2000 color difference formula ($\Delta E_{00}$).[Sharma2005, Luo2001] The clinical acceptability threshold of $\Delta E<2$ corresponds to the just-noticeable difference in human color perception.[Melgosa2001]
However, clinical skin color indices are not simple Euclidean distances. The Individual Typology Angle (ITA), widely used for Fitzpatrick skin type classification,[Chardon1991, DelBino2006] is computed as an arctangent ratio:
$\text{ITA}=\arctan\left(\frac{L^{*}-50}{b^{*}}\right)\times\frac{180}{\pi}$ (1)
This angular metric exhibits fundamentally different error sensitivity than the distance-based $\Delta E$ formula. A calibration that minimizes Euclidean color error may still harbor residual errors in specific color channels that are amplified by ratio-based clinical formulas.
Prior research has validated smartphone telemedicine for specific conditions,[Lamel2012, VanoGalvan2015] developed multi-device domain adaptation for other imaging modalities,[Chen2019, Tellez2019, Guan2021] and established RGB-based chromophore quantification methods.[Tsumura1999, Stamatas2004] However, these studies predominantly utilized datasets composed of Fitzpatrick I–II skin types,[Daneshjou2022, Groh2021] and none systematically tested whether colorimetric calibration guarantees clinical biomarker preservation.
This study addresses this critical gap. Using 43,425 images from 965 Korean subjects (Fitzpatrick III–IV), we test the hypothesis that standard colorimetric calibration ensures clinical reliability. Our investigation reveals three key findings:
  1. 1.
    Color-Clinical Decoupling: We introduce and systematically investigate “color-clinical decoupling”—the phenomenon where perceptually accurate calibration ($\Delta E<2.3$) fails to preserve clinical biomarkers. While prior work has identified ITA measurement limitations including sensitivity to illuminants,[Jung2024] this is the first systematic demonstration that standard colorimetric calibration creates false clinical confidence, with ITA showing poor agreement (ICC =0.40=0.40) despite excellent color fidelity. We provide mathematical derivation explaining this phenomenon through the differential sensitivity of ratio-based formulas to channel-specific errors.
  2. 2.
[…]
  3. 3.
    Fitzpatrick III–IV Benchmark: We provide the first large-scale multi-device benchmark specifically for the “forgotten middle” of the skin tone spectrum, addressing a critical gap in algorithmic fairness.
These findings establish that current smartphone dermatology standards create false confidence: images that “look correct” may “diagnose incorrectly.” We argue that the field must move beyond perceptual color metrics toward biomarker-aware calibration paradigms.
2 Background and Related Work
2.1 Optical Properties of Human Skin
Human skin functions as a multilayered turbid medium whose appearance is governed by the interplay of absorption and scattering within the epidermis and dermis.[Tuchin2015, Jacques2013] The two principal chromophores are melanin, concentrated in the basal layer of the epidermis, and hemoglobin, distributed within the dermal vasculature.[Anderson1981] Melanin exhibits broadband absorption across the visible spectrum, decreasing monotonically from 300 to 700 nm. Oxyhemoglobin displays characteristic absorption peaks at 542 nm and 577 nm, and blood stasis patterns further modulate perceived skin pigmentation.[Zonios2001, Stamatas2008]
The spectral reflectance of skin results from a complex light transport process involving specular reflection at the stratum corneum, volumetric scattering within the dermis, and wavelength-dependent absorption at varying depths.[Meglinski2002, Baranoski2004] RGB sensors integrate this continuous spectrum into three broad channels, irreversibly discarding the fine spectral detail required to deconvolve individual chromophore contributions—a phenomenon known as metamerism.[Nishidate2004]
2.2 Device-Dependent Color Rendering
Modern digital cameras exhibit substantial variability in color reproduction due to differences in sensor design and ISP pipelines.[Karaimer2016] Recent assessments of smartphone cameras for digital health applications have confirmed persistent color accuracy limitations across consumer devices.[Yoon2020] Three primary sources of device dependence exist: (1) spectral sensitivity functions of color filter arrays vary across manufacturers;[Hong2001, Chakrabarti2014] (2) ISP pipelines apply nonlinear transformations including gamma correction, tone mapping, and “memory color” optimization;[Karaimer2016, Cheung2004] and (3) the sRGB color space provides limited gamut and introduces quantization errors in dark tones.[Finlayson2015]
These limitations motivate the use of Color Correction Matrices (CCMs) to transform camera RGB values toward device-independent references.[Hong2001] While linear CCMs provide computational efficiency, polynomial extensions have demonstrated improved accuracy for strongly nonlinear responses.[Finlayson2015, Cheung2004]
2.3 Clinical Skin Color Indices
Quantitative dermatology employs several derived indices to characterize skin pigmentation. The Individual Typology Angle (ITA) provides an objective correlate of Fitzpatrick skin phototype.[Chardon1991, Wilkes2015, DelBino2006] The Melanin Index, computed from $L^{*}$ using logarithmic transformations, correlates with epidermal melanin concentration.[Chardon1991, Stamatas2004] The Erythema Index, derived from the $a^{*}$ channel, reflects dermal hemoglobin content.[Fullerton1996, Takiwaki1998] Critically, while device validation studies have assessed inter-device agreement for individual indices,[Yoon2020] no prior work has systematically compared how different clinical metrics—with their distinct mathematical structures—respond to identical calibration procedures. This gap is significant: a calibration optimized for one index may fail another.
2.4 The Implicit Assumption of Colorimetric Calibration
The field of tele-dermatology operates under an unstated assumption: minimizing perceptual color error ($\Delta E$) guarantees clinical validity. This assumption conflates two distinct objectives—visual fidelity and diagnostic accuracy—that may require fundamentally different optimization targets.
The CIEDE2000 formula was designed to correlate with human perceptual discrimination:[Sharma2005, Luo2001]
$\Delta E_{00}=\sqrt{\left(\frac{\Delta L^{\prime}}{k_{L}S_{L}}\right)^{2}+\left(\frac{\Delta C^{\prime}}{k_{C}S_{C}}\right)^{2}+\left(\frac{\Delta H^{\prime}}{k_{H}S_{H}}\right)^{2}+R_{T}\frac{\Delta C^{\prime}}{k_{C}S_{C}}\frac{\Delta H^{\prime}}{k_{H}S_{H}}}$ (2)
This formula applies weighting functions that suppress errors humans perceive poorly while emphasizing perceptually salient differences. Crucially, this perceptual optimization is misaligned with clinical needs.
Consider the ITA formula: $\arctan[(L^{*}-50)/b^{*}]$. This is an angular metric in the $L^{*}$-$b^{*}$ plane, fundamentally different from the distance metric of $\Delta E$. Error propagation analysis reveals:
$\delta(\text{ITA})=\frac{\partial\text{ITA}}{\partial L^{*}}\delta L^{*}+\frac{\partial\text{ITA}}{\partial b^{*}}\delta b^{*}$ (3)
where the partial derivatives are:
$\frac{\partial\text{ITA}}{\partial L^{*}}$ $=\frac{b^{*}}{(L^{*}-50)^{2}+(b^{*})^{2}}$ (4)
$\frac{\partial\text{ITA}}{\partial b^{*}}$ $=\frac{-(L^{*}-50)}{(L^{*}-50)^{2}+(b^{*})^{2}}$ (5)
For typical skin colors in our dataset ($L^{*}\approx 75$, $b^{*}\approx 15$), equation (5) yields $|\partial\text{ITA}/\partial b^{*}|\approx 0.04$ rad/unit $\approx 2.3$°/unit. A seemingly small 1-unit error in $b^{*}$ produces a 2–3° shift in ITA—sufficient to cross Fitzpatrick skin type boundaries—while contributing only $\sim$1 to the $\Delta E$ calculation. Notably, perceptibility thresholds differ across CIELAB axes: thresholds of approximately 1.0 $\Delta E$ units apply to $L^{*}$ and $a^{*}$, but $\sim$2.6 units for $b^{*}$,[LindseyWee2007] meaning $b^{*}$ errors are both less perceptible and more clinically consequential.
The $b^{*}$ channel is derived primarily from the blue sensor, which exhibits the lowest quantum efficiency among RGB channels (approximately 30% of green channel efficiency).[Ly2020] Shot noise scales as $\sigma\propto 1/\sqrt{QE}$, making blue the noisiest channel by a factor of $\sim$1.8. This noise propagates through sRGB-to-CIELAB transformation and is amplified by the ITA formula’s sensitivity to $b^{*}$.
Linear CCM minimizes the $L_{2}$ norm of color error:
$\min_{\mathbf{A},\mathbf{b}}\sum_{i=1}^{N}\left\|\mathbf{A}\cdot\text{RGB}_{i}+\mathbf{b}-\text{RGB}_{\text{ref}}^{(i)}\right\|_{2}^{2}$ (6)
This optimization reduces errors uniformly across color channels, but ITA requires disproportionate reduction in $b^{*}$ errors. The result: a calibration can achieve $\Delta E<2$ while harboring $b^{*}$ residuals that produce clinically significant ITA discrepancies.
We term this phenomenon color-clinical decoupling: the divergence between perceptual color accuracy and clinical biomarker reliability.
2.5 Dataset Bias in Dermatological AI
Dermatological AI systems suffer from severe demographic imbalances.[Adamson2018, Daneshjou2022] The ISIC Archive and other benchmarks are predominantly composed of Fitzpatrick I–II types, with limited representation of darker skin.[Kinyanjui2020, Groh2021] Furthermore, the Fitzpatrick scale itself conflates race with photosensitivity, limiting its utility for diverse populations.[Ware2020] Critically, intermediate skin tones (Fitzpatrick III–IV)—common in East Asian, Hispanic, and Mediterranean populations—represent a “forgotten middle” that is neither the focus of fairness initiatives targeting darker skin nor well-represented in existing light-skin-dominated datasets.
While underrepresentation of darker skin types in dermatological AI is well-documented,[Kinyanjui2020, Groh2021, Adamson2018] our dataset of 43,425 images from 965 Korean subjects (Fitzpatrick III–IV) represents the largest cohort specifically designed to evaluate color-clinical decoupling in this underrepresented population, providing systematic characterization of device-dependent color rendering.
3 Materials and Methods
3.1 Dataset
This study utilized the Korean Skin Condition Measurement dataset from AI Hub (aihub.or.kr), a publicly available resource for dermatological research.[AIHub2023] The dataset comprises standardized facial images captured under controlled lighting conditions.
3.1.1 Subject Demographics
A total of 965 subjects were included:
  • Age groups: $\leq$30 years (n=2,950 samples), 31–45 years (n=2,590), 46–60 years (n=2,570), >>60 years (n=1,540)
  • Gender: Female (n=4,830 samples), Male (n=4,820 samples)
  • Fitzpatrick skin types: Predominantly Types III–IV
3.1.2 Imaging Devices
Three device categories were employed:
  1. 1.
    DSLR (Reference): Professional-grade camera under controlled dark room lighting (33,774 samples, 7 angles)
  2. 2.
    Tablet: Consumer-grade tablet camera (4,825 samples, 3 angles)
  3. 3.
    Smartphone: Consumer-grade smartphone camera (4,825 samples, 3 angles)
3.1.3 Facial Regions
Five standardized facial regions were analyzed: forehead, left cheek, right cheek, chin, and glabella.
3.2 Color Space Transformation
RGB values were transformed to CIELAB using D65 illuminant:
$\text{RGB}\xrightarrow{\text{Gamma}}\text{Linear RGB}\xrightarrow{\text{D65}}\text{XYZ}\xrightarrow{\text{CIE}}L^{*}a^{*}b^{*}$ (7)
3.3 Color Correction Matrix
Linear CCM normalization was performed using ordinary least squares:
$\begin{bmatrix}R_{\text{corr}}\\ G_{\text{corr}}\\ B_{\text{corr}}\end{bmatrix}=\mathbf{A}\begin{bmatrix}R_{\text{src}}\\ G_{\text{src}}\\ B_{\text{src}}\end{bmatrix}+\mathbf{b}$ (8)
where $\mathbf{A}\in\mathbb{R}^{3\times 3}$ and $\mathbf{b}\in\mathbb{R}^{3}$. Validation employed 5-fold cross-validation with 80/20 train-test split.
3.4 Clinical Index Calculations
3.4.1 Melanin Index
$\text{MI}=100\times\log_{10}\left(\frac{100}{L^{*}}\right)$ (9)
3.4.2 Erythema Index
$\text{EI}=a^{*}-0.5\times L^{*}$ (10)
3.4.3 Individual Typology Angle
$\text{ITA}=\arctan\left(\frac{L^{*}-50}{b^{*}}\right)\times\frac{180}{\pi}$ (11)
3.5 Statistical Analysis
Inter-device agreement was assessed using ICC(3,1) for absolute agreement.[Shrout1979, Koo2016] ICC interpretation: <<0.50 (poor), 0.50–0.75 (moderate), 0.75–0.90 (good), >>0.90 (excellent).[Portney2015]
Bland-Altman analysis characterized device agreement with 95% limits of agreement.[Bland1986]
ANOVA assessed factor influences on $\Delta E$, with effect sizes quantified using eta-squared ($\eta^{2}$): 0.01 (small), 0.06 (medium), 0.14 (large).[Cohen1988]
All analyses used Python 3.8 with scipy, pingouin, and statsmodels. Significance: $\alpha=0.05$ with Bonferroni correction.
4 Results
4.1 Raw Device Performance: Clinical Unusability of Uncalibrated Images
Table 1 summarizes device characteristics. Consumer devices exhibited systematic underexposure relative to DSLR ($L^{*}=72.7$$77.6$ vs. $81.3$) and compressed $b^{*}$ values (8.968.9610.1210.12 vs. $17.59$).
Table 1: Device characterization summary.
Parameter DSLR Tablet Smartphone
N samples 33,774 4,825 4,825
Mean $L^{*}$ 81.35 72.73 77.56
Mean $a^{*}$ 7.95 6.08 5.68
Mean $b^{*}$ 17.59 10.12 8.96
SNR (LAB) 37.5 dB
[…]
Table 2: Inter-device CIEDE2000 color differences.
Device Pair Mean Std Median 95th %ile
DSLR vs Tablet 8.26 2.70 7.57 13.71
DSLR vs Smartphone 6.99 1.84 6.77 10.37
Tablet vs Smartphone 4.47 2.39 4.13 8.97
This finding empirically demonstrates that raw smartphone images are clinically unusable for quantitative skin analysis without active calibration.
Figure 1: RGB and CIELAB distributions across devices showing systematic shifts in color rendering.
Figure 2: Bland-Altman plots for $L^{*}$, $a^{*}$, $b^{*}$. Largest bias in $b^{*}$ channel (DSLR vs smartphone: 8.30).
4.2 CCM Normalization: Colorimetric Success
Linear CCM substantially reduced color differences (Table 3): 77.4% improvement for tablet ($\Delta E$: 8.33 $\rightarrow$ 1.88) and 67.4% for smartphone (7.02 $\rightarrow$ 2.29). Post-normalization median values approached clinical acceptability (1.66 and 2.09 respectively).
Table 3: CCM normalization performance (5-fold cross-validation).
Transformation $\Delta E$ Before $\Delta E$ After Improvement
Tablet $\rightarrow$ DSLR $8.33\pm 2.76$ $1.88\pm 1.14$ 77.4%
Smartphone $\rightarrow$ DSLR $7.02\pm 1.78$ $2.29\pm 1.22$ 67.4%
Figure 3: $\Delta E$ distributions before and after CCM calibration.
4.3 The Decoupling: Clinical Index Failure Despite Colorimetric Success
Despite effective colorimetric normalization, clinical index agreement varied dramatically (Table 4). While Melanin Index achieved good agreement (ICC =0.767=0.767), ITA showed poor agreement (ICC =0.402=0.402). This represents the central finding: perceptual color accuracy does not guarantee clinical biomarker reliability.
Table 4: Intraclass Correlation Coefficients for clinical indices and CIELAB channels.
Index/Channel ICC(3,1) 95% CI Interpretation
Melanin Index 0.767 [0.76, 0.78] Good
Erythema Index 0.652 [0.64, 0.66] Moderate
ITA 0.402 [0.38, 0.42] Poor
$L^{*}$ 0.774 [0.76, 0.78] Good
$a^{*}$ 0.550 [0.53, 0.57] Moderate
$b^{*}$ 0.425 [0.41, 0.44] Poor
The pattern in Table 4 reveals the mechanism: $b^{*}$ channel agreement (ICC =0.425=0.425) is substantially lower than $L^{*}$ (ICC =0.774=0.774). Since ITA depends critically on $b^{*}$ through its ratio formula, the poor $b^{*}$ agreement propagates directly to ITA. In contrast, Melanin Index depends only on $L^{*}$ and thus inherits the good $L^{*}$ agreement.
Figure 4: Clinical index distributions across devices after CCM normalization.
Figure 5: Correlation plots for clinical indices. Note the scatter in ITA despite good Melanin Index correlation.
4.4 Sources of Variance: Implications for Calibration Protocol Design
ANOVA revealed a striking hierarchy of variance sources (Table 5). Facial anatomical region accounted for 25.2% of variance ($\eta^{2}=0.252$)—3.6 times greater than device hardware effects ($\eta^{2}=0.070$).
Table 5: ANOVA: Factors influencing inter-device $\Delta E$ (pre-normalization).
Factor F-statistic p-value $\eta^{2}$ Effect Size
Facial Region 813.89 <<0.001 0.252 Large
Source Device 728.56 <<0.001 0.070 Medium
Gender 72.84 <<0.001 0.007 Small
Skin Type 2.77 0.017 0.001 Small
Age Group 4.47 0.004 0.001 Small
The chin region exhibited the largest color differences (mean $\Delta E=9.63$), while the forehead showed better agreement (mean $\Delta E=6.21$). This finding challenges the prevailing single-patch calibration paradigm: a correction matrix derived from cheek measurements may fail when applied to the forehead or chin.
Figure 6: $\Delta E$ distributions by facial region revealing substantial anatomical variation.
Figure 7: Stratified $\Delta E$ analysis: (A) Age, (B) Gender, (C) Skin type, (D) Facial region.
5 Discussion
5.1 The Color-Clinical Decoupling Phenomenon
The central finding of this study is the non-linear decoupling between colorimetric accuracy and clinical biomarker reliability. Despite achieving near-clinical color fidelity ($\Delta E<2.3$), the ITA index exhibited poor inter-device agreement (ICC =0.40=0.40). This finding challenges the fundamental assumption underlying smartphone dermatology: that “looking correct” guarantees “diagnosing correctly.”
The mathematical origins of this decoupling merit analysis. The CIEDE2000 formula measures a weighted Euclidean distance in color space, optimized for human perceptual discrimination. Linear CCM minimizes the $L_{2}$ norm of RGB residuals, distributing error reduction uniformly across channels. However, the ITA formula is not a distance—it is an arctangent ratio that exhibits singular behavior as $b^{*}$ approaches zero and nonlinear amplification of $b^{*}$ errors.
From equations (4) and (5), for our dataset’s typical skin values ($L^{*}\approx 75$, $b^{*}\approx 15$):
$\left|\frac{\partial\text{ITA}}{\partial b^{*}}\right|$ $=\frac{|75-50|}{(75-50)^{2}+15^{2}}=\frac{25}{850}\approx 0.029\text{ rad/unit}\approx 1.7\text{\textdegree/unit}$ (12)
A 1-unit residual in $b^{*}$—easily masked within a $\Delta E<2$ calibration—produces a 1.7° shift in ITA. Since Fitzpatrick skin type boundaries span approximately 10° in ITA,[DelBino2006] a 3-unit $b^{*}$ error can cause skin type misclassification.
The blue sensor’s low quantum efficiency ($\sim$30% of green) produces elevated shot noise that propagates directly to $b^{*}$. Linear CCM, optimized for overall color fidelity, cannot selectively suppress this critical error source. The result is “color-clinical decoupling”: colorimetric success masking clinical failure.
In contrast, the Melanin Index formula $\text{MI}=100\times\log_{10}(100/L^{*})$ depends solely on $L^{*}$, which benefits from higher signal-to-noise in the green channel. This explains why Melanin Index achieved good agreement (ICC =0.77=0.77) while ITA failed—the two indices sample different axes of the colorimetric error ellipsoid, with ITA maximally sensitive to the noisiest component.
This finding has immediate implications for tele-dermatology: standard colorimetric calibration, the current industry norm, is insufficient for clinical applications requiring ITA or similar ratio-based biomarkers.
5.2 Why Single-Patch Calibration Fails
The color-clinical decoupling is further exacerbated by an overlooked calibration assumption: that facial skin is a uniform surface. While regional variation in facial skin color is well-documented—studies have shown significant differences in color indices between forehead, cheek, and chin regions[Pan2020, Wang2018]—prior work has not quantified regional variance as a proportion of total variance nor compared it against device variance. Our ANOVA analysis reveals this assumption to be fundamentally flawed: facial region explains 25.2% of color variance, exceeding device effects (7.0%) by a factor of 3.6—the first such quantitative decomposition.
This “anatomical dominance” arises from multiple physiological sources. The T-zone (forehead and glabella) exhibits higher sebaceous gland density than cheek regions, resulting in elevated sebum secretion that alters surface refractive index.[Man2009, Ohshima2013] Sebum forms a thin lipid film that increases specular reflection and modifies the balance between surface gloss and diffuse reflectance.[Baranoski2004] Under directional illumination, this regional variation produces angle-dependent color shifts.
Facial skin thickness varies substantially: the periorbital region is among the thinnest (0.5–1.0 mm), while chin and forehead are thicker (1.5–2.0 mm).[Nouveau2016, Pioro2021] Since skin color arises from chromophore absorption at specific depths, thickness variations alter the optical path length and relative contributions of melanin versus hemoglobin.
The chin region, exhibiting the largest color differences (mean $\Delta E=9.63$), presents a particularly challenging case due to pronounced surface curvature creating geometric variations in illumination.
This finding invalidates single-patch calibration protocols: a CCM derived from cheek measurements may fail when applied to forehead or chin. Future calibration must be region-aware, with region-specific corrections or spatially-varying calibration fields based on facial landmark detection.
5.3 Implications for Tele-Dermatology Standards
Our findings have critical implications for smartphone dermatology. First, the 0% raw acceptability rate proves that uncalibrated consumer devices are clinically unusable for quantitative analysis. The widespread practice of using raw smartphone images for AI training likely introduces device-specific biases that obscure pathological signals.[Adamson2018, Daneshjou2022]
Second, even with calibration, the color-clinical decoupling renders current standards inadequate for ITA-dependent applications, including Fitzpatrick skin type classification and melanoma risk assessment. Clinical applications must either: (a) avoid ratio-based indices, (b) develop biomarker-aware calibration that explicitly minimizes ITA error, or (c) employ uncertainty quantification to flag unreliable measurements.
Third, the statistical significance of gender effects in our ANOVA ($p<0.001$) suggests that camera ISPs may exhibit differential behavior across demographics. If “universal” CCMs are derived from specific populations, they may embed demographic biases—a fairness concern for global health applications.[Kinyanjui2020, Groh2021]
5.4 Limitations
Several limitations should be acknowledged. First, we identified the decoupling phenomenon but did not implement non-linear or biomarker-aware corrections; such approaches remain for future investigation. Second, our dataset was restricted to healthy skin; pathological tissue may exhibit different spectral properties. Third, the population was limited to Korean subjects (Fitzpatrick III–IV); generalization to types I–II and V–VI requires additional study. Fourth, controlled lighting conditions may not represent real-world acquisition environments.
5.5 Future Directions
The color-clinical decoupling motivates several research directions:
Biomarker-Specific Loss Functions. Composite losses combining perceptual ($\mathcal{L}_{\Delta E}$) and clinical ($\mathcal{L}_{\text{ITA}}$) terms could optimize simultaneously for visual fidelity and diagnostic accuracy, explicitly penalizing $b^{*}$ errors.
Region-Aware Calibration. Modern facial landmark detection enables region-specific CCMs or spatially-varying correction fields accounting for local optical properties.
Spectral Reconstruction. Methods recovering hyperspectral reflectance from RGB inputs[Arad2016, Shi2018] could compute clinical indices from chromophore spectra rather than device-dependent colors, potentially resolving the decoupling.
Physics-Informed Neural Networks. PINNs embedding radiative transfer constraints[Raissi2019] could enforce biophysically plausible chromophore concentrations.
Multi-Ethnic Benchmark Extension. Extension to Fitzpatrick I–VI and the Monk Skin Tone scale[Monk2023] is essential for equitable algorithm development.
6 Conclusion
This study establishes that standard colorimetric calibration in smartphone dermatology creates false confidence: perceptually accurate images can yield unreliable clinical biomarkers. Using 43,425 images from 965 Korean subjects across DSLR, tablet, and smartphone devices, we demonstrate a fundamental “color-clinical decoupling” phenomenon.
Our key findings are:
  1. 1.
    Raw images are clinically unusable: 0% of uncalibrated consumer device images met the $\Delta E<2$ clinical threshold, with mean errors of 6.99–8.26.
  2. 2.
[…]
Additionally, this work provides the first large-scale multi-device benchmark specifically for Fitzpatrick III–IV skin tones, addressing the “forgotten middle” in algorithmic fairness research.
The color-clinical decoupling arises from a mathematical mismatch: $\Delta E$ is a distance metric while ITA is an angular ratio disproportionately sensitive to $b^{*}$ errors—the noisiest color channel. Linear CCM, minimizing overall color distance, cannot selectively suppress this critical error source.
These findings argue that tele-dermatology must move beyond perceptual color metrics toward biomarker-aware calibration paradigms. Current smartphone apps using ITA for skin type classification may be producing unreliable results even when images “look correct.” Future calibration protocols must be region-aware, biomarker-specific, and validated on diverse populations to ensure both accuracy and algorithmic fairness.
Disclosures
The authors declare no conflicts of interest.
Code, Data, and Materials Availability
The Korean Skin Condition Measurement dataset is available through AI Hub (aihub.or.kr) under restricted research access. Analysis code will be available on GitHub upon publication.

### v2

@@
 Objective: To investigate whether standard colorimetric calibration translates to clinical biomarker reliability, and to identify the dominant sources of variance in cross-device skin color measurement.
-Methods: We analyzed 43,425 images from 965 Korean subjects (Fitzpatrick III–IV) across DSLR, tablet, and smartphone devices using the AI Hub Korean Skin Condition Measurement dataset.
+Methods: We analyzed a total dataset of 43,424 images from 965 Korean subjects (Fitzpatrick III–IV) using the AI Hub Korean Skin Condition Measurement dataset.
+Cross-device comparisons utilized 14,475 matched front-facing images (4,825 per device) acquired across DSLR, tablet, and smartphone devices.
 Linear Color Correction Matrix (CCM) normalization—the industry standard—was applied to align consumer devices to a DSLR reference.
@@
 Skin functions as a multilayered turbid medium, with color determined by the absorption spectra of melanin in the epidermis and hemoglobin in the dermal vasculature.[Jacques2013, Anderson1981] When this continuous spectral information is sampled by RGB sensors, the fine spectral detail required to separate chromophore contributions is irreversibly lost.[Zonios2001, Nishidate2004] Device-specific variations compound this limitation: different sensors produce inconsistent RGB values for identical skin samples due to variations in spectral sensitivities and image signal processing (ISP) pipelines.[Finlayson2015, Karaimer2016]
-The standard solution—linear Color Correction Matrix (CCM) normalization—transforms device RGB values toward a reference standard by minimizing the Euclidean distance in color space.[Hong2001, Cheung2004] This approach optimizes for perceptual color fidelity, as measured by the CIEDE2000 color difference formula ($\Delta E_{00}$).[Sharma2005, Luo2001] The clinical acceptability threshold of $\Delta E<2$ corresponds to the just-noticeable difference in human color perception.[Melgosa2001]
+The standard solution—linear Color Correction Matrix (CCM) normalization—transforms device RGB values toward a reference standard by minimizing the Euclidean distance in color space.[Hong2001, Cheung2004] This approach optimizes for perceptual color fidelity, as measured by the CIEDE2000 color difference formula ($\Delta E_{00}$).[Sharma2005, Luo2001] The clinical acceptability threshold of $\Delta E<2$ corresponds to the just-noticeable difference in human color perception.[Melgosa1997]
 However, clinical skin color indices are not simple Euclidean distances.
@@
 A calibration that minimizes Euclidean color error may still harbor residual errors in specific color channels that are amplified by ratio-based clinical formulas.
-Prior research has validated smartphone telemedicine for specific conditions,[Lamel2012, VanoGalvan2015] developed multi-device domain adaptation for other imaging modalities,[Chen2019, Tellez2019, Guan2021] and established RGB-based chromophore quantification methods.[Tsumura1999, Stamatas2004] However, these studies predominantly utilized datasets composed of Fitzpatrick I–II skin types,[Daneshjou2022, Groh2021] and none systematically tested whether colorimetric calibration guarantees clinical biomarker preservation.
+Prior research has validated smartphone telemedicine for specific conditions,[Lamel2012, VanoGalvan2011] developed multi-device domain adaptation for other imaging modalities,[Chen2019, Tellez2019, Guan2021] and established RGB-based chromophore quantification methods.[Tsumura2003, Stamatas2004] However, these studies predominantly utilized datasets composed of Fitzpatrick I–II skin types,[Daneshjou2022, Groh2021] and none systematically tested whether colorimetric calibration guarantees clinical biomarker preservation.
 This study addresses this critical gap.
-Using 43,425 images from 965 Korean subjects (Fitzpatrick III–IV), we test the hypothesis that standard colorimetric calibration ensures clinical reliability.
+Using 43,424 images from 965 Korean subjects (Fitzpatrick III–IV), we test the hypothesis that standard colorimetric calibration ensures clinical reliability.
 Our investigation reveals three key findings:
@@
     Color-Clinical Decoupling: We introduce and systematically investigate “color-clinical decoupling”—the phenomenon where perceptually accurate calibration ($\Delta E<2.3$) fails to preserve clinical biomarkers.
-While prior work has identified ITA measurement limitations including sensitivity to illuminants,[Jung2024] this is the first systematic demonstration that standard colorimetric calibration creates false clinical confidence, with ITA showing poor agreement (ICC =0.40=0.40) despite excellent color fidelity.
+While prior work has identified ITA measurement limitations including sensitivity to illuminants,[Jung2024] this is the first systematic demonstration that standard colorimetric calibration creates false clinical confidence, with ITA showing poor agreement (ICC =0.46=0.46) despite excellent color fidelity.
 We provide mathematical derivation explaining this phenomenon through the differential sensitivity of ratio-based formulas to channel-specific errors.
@@
 2.1 Optical Properties of Human Skin
-Human skin functions as a multilayered turbid medium whose appearance is governed by the interplay of absorption and scattering within the epidermis and dermis.[Tuchin2015, Jacques2013] The two principal chromophores are melanin, concentrated in the basal layer of the epidermis, and hemoglobin, distributed within the dermal vasculature.[Anderson1981] Melanin exhibits broadband absorption across the visible spectrum, decreasing monotonically from 300 to 700 nm.
-Oxyhemoglobin displays characteristic absorption peaks at 542 nm and 577 nm, and blood stasis patterns further modulate perceived skin pigmentation.[Zonios2001, Stamatas2008]
-The spectral reflectance of skin results from a complex light transport process involving specular reflection at the stratum corneum, volumetric scattering within the dermis, and wavelength-dependent absorption at varying depths.[Meglinski2002, Baranoski2004] RGB sensors integrate this continuous spectrum into three broad channels, irreversibly discarding the fine spectral detail required to deconvolve individual chromophore contributions—a phenomenon known as metamerism.[Nishidate2004]
+Human skin functions as a multilayered turbid medium whose appearance is governed by the interplay of absorption and scattering within the epidermis and dermis.[Tuchin2000, Jacques2013] The two principal chromophores are melanin, concentrated in the basal layer of the epidermis, and hemoglobin, distributed within the dermal vasculature.[Anderson1981] Melanin exhibits broadband absorption across the visible spectrum, decreasing monotonically from 300 to 700 nm.
+Oxyhemoglobin displays characteristic absorption peaks at 542 nm and 577 nm, and blood stasis patterns further modulate perceived skin pigmentation.[Zonios2001, Stamatas2004b]
+The spectral reflectance of skin results from a complex light transport process involving specular reflection at the stratum corneum, volumetric scattering within the dermis, and wavelength-dependent absorption at varying depths.[Meglinski2002, Baranoski2004] RGB sensors integrate this continuous spectrum into three broad channels, irreversibly discarding the fine spectral detail required to deconvolve individual chromophore contributions—a limitation arising from metamerism.
 2.2 Device-Dependent Color Rendering
-Modern digital cameras exhibit substantial variability in color reproduction due to differences in sensor design and ISP pipelines.[Karaimer2016] Recent assessments of smartphone cameras for digital health applications have confirmed persistent color accuracy limitations across consumer devices.[Yoon2020] Three primary sources of device dependence exist: (1) spectral sensitivity functions of color filter arrays vary across manufacturers;[Hong2001, Chakrabarti2014] (2) ISP pipelines apply nonlinear transformations including gamma correction, tone mapping, and “memory color” optimization;[Karaimer2016, Cheung2004] and (3) the sRGB color space provides limited gamut and introduces quantization errors in dark tones.[Finlayson2015]
+Modern digital cameras exhibit substantial variability in color reproduction due to differences in sensor design and ISP pipelines.[Karaimer2016] Three primary sources of device dependence exist: (1) spectral sensitivity functions of color filter arrays vary across manufacturers;[Hong2001, Chakrabarti2014] (2) ISP pipelines apply nonlinear transformations including gamma correction, tone mapping, and “memory color” optimization;[Karaimer2016, Cheung2004] and (3) the sRGB color space provides limited gamut and introduces quantization errors in dark tones.
 These limitations motivate the use of Color Correction Matrices (CCMs) to transform camera RGB values toward device-independent references.[Hong2001] While linear CCMs provide computational efficiency, polynomial extensions have demonstrated improved accuracy for strongly nonlinear responses.[Finlayson2015, Cheung2004]
@@
 Quantitative dermatology employs several derived indices to characterize skin pigmentation.
-The Individual Typology Angle (ITA) provides an objective correlate of Fitzpatrick skin phototype.[Chardon1991, Wilkes2015, DelBino2006] The Melanin Index, computed from $L^{*}$ using logarithmic transformations, correlates with epidermal melanin concentration.[Chardon1991, Stamatas2004] The Erythema Index, derived from the $a^{*}$ channel, reflects dermal hemoglobin content.[Fullerton1996, Takiwaki1998] Critically, while device validation studies have assessed inter-device agreement for individual indices,[Yoon2020] no prior work has systematically compared how different clinical metrics—with their distinct mathematical structures—respond to identical calibration procedures.
+The Individual Typology Angle (ITA) provides an objective correlate of Fitzpatrick skin phototype.[Chardon1991, Wilkes2015, DelBino2006] The Melanin Index, computed from $L^{*}$ using logarithmic transformations, correlates with epidermal melanin concentration.[Chardon1991, Stamatas2004] The Erythema Index, derived from the $a^{*}$ channel, reflects dermal hemoglobin content.[Fullerton1996, Takiwaki1998] Critically, while device validation studies have assessed inter-device agreement for individual indices, no prior work has systematically compared how different clinical metrics—with their distinct mathematical structures—respond to identical calibration procedures.
 This gap is significant: a calibration optimized for one index may fail another.
@@
 $\frac{\partial\text{ITA}}{\partial b^{*}}$ $=\frac{-(L^{*}-50)}{(L^{*}-50)^{2}+(b^{*})^{2}}$ (5)
-For typical skin colors in our dataset ($L^{*}\approx 75$, $b^{*}\approx 15$), equation (5) yields $|\partial\text{ITA}/\partial b^{*}|\approx 0.04$ rad/unit $\approx 2.3$°/unit.
-A seemingly small 1-unit error in $b^{*}$ produces a 2–3° shift in ITA—sufficient to cross Fitzpatrick skin type boundaries—while contributing only $\sim$1 to the $\Delta E$ calculation.
-Notably, perceptibility thresholds differ across CIELAB axes: thresholds of approximately 1.0 $\Delta E$ units apply to $L^{*}$ and $a^{*}$, but $\sim$2.6 units for $b^{*}$,[LindseyWee2007] meaning $b^{*}$ errors are both less perceptible and more clinically consequential.
-The $b^{*}$ channel is derived primarily from the blue sensor, which exhibits the lowest quantum efficiency among RGB channels (approximately 30% of green channel efficiency).[Ly2020] Shot noise scales as $\sigma\propto 1/\sqrt{QE}$, making blue the noisiest channel by a factor of $\sim$1.8.
+For typical skin colors in our dataset ($L^{*}\approx 75$, $b^{*}\approx 15$), equation (5) yields $|\partial\text{ITA}/\partial b^{*}|\approx 0.03$ rad/unit $\approx 1.7$°/unit.
+A seemingly small 1-unit error in $b^{*}$ produces a 1.7° shift in ITA—sufficient to cross Fitzpatrick skin type boundaries—while contributing only $\sim$1 to the $\Delta E$ calculation.
+Notably, perceptibility thresholds differ across CIELAB axes, with the $b^{*}$ (blue-yellow) axis generally exhibiting higher tolerance thresholds than $L^{*}$ or $a^{*}$, meaning $b^{*}$ errors are both less perceptible and more clinically consequential.
+The $b^{*}$ channel is derived primarily from the blue sensor.
+Due to the shallow absorption depth of blue photons in silicon and the spectral deficiency of common indoor illuminants, the effective signal-to-noise ratio of the blue channel is significantly lower than that of the green channel.[Nakamura2005, Fossum1997] Shot noise scales as $\sigma\propto 1/\sqrt{QE}$, making blue the noisiest channel.
 This noise propagates through sRGB-to-CIELAB transformation and is amplified by the ITA formula’s sensitivity to $b^{*}$.
@@
 Dermatological AI systems suffer from severe demographic imbalances.[Adamson2018, Daneshjou2022] The ISIC Archive and other benchmarks are predominantly composed of Fitzpatrick I–II types, with limited representation of darker skin.[Kinyanjui2020, Groh2021] Furthermore, the Fitzpatrick scale itself conflates race with photosensitivity, limiting its utility for diverse populations.[Ware2020] Critically, intermediate skin tones (Fitzpatrick III–IV)—common in East Asian, Hispanic, and Mediterranean populations—represent a “forgotten middle” that is neither the focus of fairness initiatives targeting darker skin nor well-represented in existing light-skin-dominated datasets.
-While underrepresentation of darker skin types in dermatological AI is well-documented,[Kinyanjui2020, Groh2021, Adamson2018] our dataset of 43,425 images from 965 Korean subjects (Fitzpatrick III–IV) represents the largest cohort specifically designed to evaluate color-clinical decoupling in this underrepresented population, providing systematic characterization of device-dependent color rendering.
+While underrepresentation of darker skin types in dermatological AI is well-documented,[Kinyanjui2020, Groh2021, Adamson2018] our dataset of 43,424 images from 965 Korean subjects (Fitzpatrick III–IV) represents the largest cohort specifically designed to evaluate color-clinical decoupling in this underrepresented population, providing systematic characterization of device-dependent color rendering.
 3 Materials and Methods
@@
   1. 1.
-    DSLR (Reference): Professional-grade camera under controlled dark room lighting (33,774 samples, 7 angles)
+    DSLR (Reference): Professional-grade camera under controlled dark room lighting, capturing 7 viewing angles (front, front-top, front-bottom, left 15°, left 30°, right 15°, right 30°)
   2. 2.
-    Tablet: Consumer-grade tablet camera (4,825 samples, 3 angles)
+    Tablet: Consumer-grade tablet camera, capturing 3 viewing angles (front, left, right)
   3. 3.
-    Smartphone: Consumer-grade smartphone camera (4,825 samples, 3 angles)
+    Smartphone: Consumer-grade smartphone camera, capturing 3 viewing angles (front, left, right)
+The analyzed dataset comprises 43,424 samples: DSLR (33,774 samples across 7 angles), tablet (4,825 front-facing samples), and smartphone (4,825 front-facing samples).
+One DSLR sample was excluded due to invalid bounding box coordinates that prevented reliable color extraction.
+Although the source dataset includes 3 angles for consumer devices (totaling 14,475 samples per device), only the front-facing imagery was included in this study to maintain focus on cross-device consistency.
+For cross-device agreement analysis, these 14,475 matched front-facing samples (4,825 per device) ensured geometric alignment across all three platforms.
 3.1.3 Facial Regions
@@
 $81.3$) and compressed $b^{*}$ values (8.968.9610.1210.12 vs.
-$17.59$).
+$17.59$), as visualized in Figure 1 and Figure 2.
 Table 1: Device characterization summary.
@@
 This finding empirically demonstrates that raw smartphone images are clinically unusable for quantitative skin analysis without active calibration.
-Figure 1: RGB and CIELAB distributions across devices showing systematic shifts in color rendering.
-Figure 2: Bland-Altman plots for $L^{*}$, $a^{*}$, $b^{*}$. Largest bias in $b^{*}$ channel (DSLR vs smartphone: 8.30).
+Figure 1: RGB and CIELAB color distributions across three imaging devices (DSLR, tablet, smartphone) for Korean skin samples (total n=43,424; cross-device comparison subset n=14,475).
+(A) RGB channel distributions showing device-specific color rendering characteristics.
+(B) CIELAB $L^{*}a^{*}b^{*}$ distributions demonstrating lightness and chromaticity variations across devices.
+Figure 2: Signal-to-noise ratio (SNR) comparison across devices calculated from repeated measurements at multiple capture angles.
+Higher SNR values indicate better measurement precision and lower random noise.
+Error bars represent standard deviation across facial regions (forehead, left cheek, right cheek, chin, glabella).
+Figure 3: Bland-Altman plots comparing $L^{*}$, $a^{*}$, and $b^{*}$ measurements between consumer devices (tablet, smartphone) and reference DSLR.
+Solid horizontal lines indicate mean bias; dashed lines show 95% limits of agreement ($\pm$1.96 SD).
+Figure 4: CIEDE2000 color difference ($\Delta E$) distributions by facial region and source device.
+The vertical dashed line indicates the clinically perceptible threshold ($\Delta E=2.0$).
+The chin region exhibits the largest color differences, while cheek regions show better inter-device agreement.
 4.2 CCM Normalization: Colorimetric Success
-Linear CCM substantially reduced color differences (Table 3): 77.4% improvement for tablet ($\Delta E$: 8.33 $\rightarrow$ 1.88) and 67.4% for smartphone (7.02 $\rightarrow$ 2.29).
+Linear CCM substantially reduced color differences (Table 3 and Figure 5): 77.4% improvement for tablet ($\Delta E$: 8.33 $\rightarrow$ 1.88) and 67.4% for smartphone (7.02 $\rightarrow$ 2.29).
 Post-normalization median values approached clinical acceptability (1.66 and 2.09 respectively).
@@
 Smartphone $\rightarrow$ DSLR $7.02\pm 1.78$ $2.29\pm 1.22$ 67.4%
-Figure 3: $\Delta E$ distributions before and after CCM calibration.
+Figure 5: Color correction matrix (CCM) normalization results showing $\Delta E$ distributions before and after calibration.
+Post-normalization $\Delta E$ distributions demonstrate significant improvement, with linear CCM trained using 5-fold cross-validation.
 4.3 The Decoupling: Clinical Index Failure Despite Colorimetric Success
-Despite effective colorimetric normalization, clinical index agreement varied dramatically (Table 4).
-While Melanin Index achieved good agreement (ICC =0.767=0.767), ITA showed poor agreement (ICC =0.402=0.402).
+Despite effective colorimetric normalization, clinical index agreement varied dramatically (Table 4, Figure 6, and Figure 7).
+While Melanin Index achieved good agreement (ICC =0.767=0.767), ITA showed poor agreement (ICC =0.455=0.455).
 This represents the central finding: perceptual color accuracy does not guarantee clinical biomarker reliability.
@@
 Erythema Index 0.652 [0.64, 0.66] Moderate
-ITA 0.402 [0.38, 0.42] Poor
+ITA 0.455 [0.44, 0.47] Poor
 $L^{*}$ 0.774 [0.76, 0.78] Good
@@
 In contrast, Melanin Index depends only on $L^{*}$ and thus inherits the good $L^{*}$ agreement.
-Figure 4: Clinical index distributions across devices after CCM normalization.
-Figure 5: Correlation plots for clinical indices. Note the scatter in ITA despite good Melanin Index correlation.
+Figure 6: Comparison of clinical skin color indices across devices after CCM normalization.
+(A) Melanin Index distributions.
+(B) Erythema Index distributions.
+(C) Individual Typology Angle (ITA) for Fitzpatrick skin type classification.
+Figure 7: Pearson correlation plots comparing clinical indices between consumer devices and DSLR reference.
+Each point represents a matched sample (same subject, same facial region).
+Diagonal dashed line indicates perfect agreement ($r=1.0$).
 4.4 Sources of Variance: Implications for Calibration Protocol Design
-ANOVA revealed a striking hierarchy of variance sources (Table 5).
+ANOVA revealed a striking hierarchy of variance sources (Table 5 and Figure 8).
 Facial anatomical region accounted for 25.2% of variance ($\eta^{2}=0.252$)—3.6 times greater than device hardware effects ($\eta^{2}=0.070$).
@@
 This finding challenges the prevailing single-patch calibration paradigm: a correction matrix derived from cheek measurements may fail when applied to the forehead or chin.
-Figure 6: $\Delta E$ distributions by facial region revealing substantial anatomical variation.
-Figure 7: Stratified $\Delta E$ analysis: (A) Age, (B) Gender, (C) Skin type, (D) Facial region.
+Figure 8: Stratified analysis of pre-normalization color differences ($\Delta E$) across demographic factors.
+(A) Age groups.
+(B) Gender.
+(C) Fitzpatrick skin type.
+(D) Facial region.
+One-way ANOVA revealed significant effects for all factors, with facial region showing the largest effect size.
 5 Discussion
@@
 The central finding of this study is the non-linear decoupling between colorimetric accuracy and clinical biomarker reliability.
-Despite achieving near-clinical color fidelity ($\Delta E<2.3$), the ITA index exhibited poor inter-device agreement (ICC =0.40=0.40).
+Despite achieving near-clinical color fidelity ($\Delta E<2.3$), the ITA index exhibited poor inter-device agreement (ICC =0.46=0.46).
 This finding challenges the fundamental assumption underlying smartphone dermatology: that “looking correct” guarantees “diagnosing correctly.”
@@
 Since Fitzpatrick skin type boundaries span approximately 10° in ITA,[DelBino2006] a 3-unit $b^{*}$ error can cause skin type misclassification.
-The blue sensor’s low quantum efficiency ($\sim$30% of green) produces elevated shot noise that propagates directly to $b^{*}$.
+Critically, this sensitivity is phenotype-dependent.
+The ITA sensitivity formula $|\partial\text{ITA}/\partial b^{*}|=|L^{*}-50|/[(L^{*}-50)^{2}+(b^{*})^{2}]$ reveals that darker skin types—characterized by lower $L^{*}$ values closer to 50 and lower $b^{*}$ values—experience dramatically amplified error sensitivity.
+For Type V skin ($L^{*}\approx 60$, $b^{*}\approx 10$), the sensitivity increases to $\sim$2.9°/unit; for Type VI skin ($L^{*}\approx 40$, $b^{*}\approx 5$), it reaches $\sim$4.6°/unit.
+A 2-unit calibration residual in $b^{*}$—perceptually negligible—produces a 3.4° shift in ITA for light skin but a 9.2° shift for dark skin.
+This represents a “metrological bias”: current calibration protocols, by optimizing for average $\Delta E$, do not suppress $b^{*}$ errors with the aggression required for accurate assessment of darker skin tones.
+Table 6 summarizes this phenotype-dependent vulnerability.
+The blue sensor’s inherently low quantum efficiency produces elevated shot noise that propagates directly to $b^{*}$.
 Linear CCM, optimized for overall color fidelity, cannot selectively suppress this critical error source.
@@
 This finding has immediate implications for tele-dermatology: standard colorimetric calibration, the current industry norm, is insufficient for clinical applications requiring ITA or similar ratio-based biomarkers.
+Table 6: Phenotype-dependent ITA sensitivity to $b^{*}$ channel errors.
+Sensitivity calculated from $|\partial\text{ITA}/\partial b^{*}|=|L^{*}-50|/[(L^{*}-50)^{2}+(b^{*})^{2}]$.
+The “2-unit Error Impact” shows the ITA shift from a typical calibration residual, demonstrating disproportionate effects on darker skin types.
+Skin Type $L^{*}$ $b^{*}$ Sensitivity (°/unit) 2-unit Error Impact
+Type II/III (Light) 75 15 1.68 3.4°
+Type IV (Intermediate) 65 15 1.92 3.8°
+Type V (Brown) 60 10 2.86 5.7°
+Type VI (Dark) 40 5 4.58 9.2°
 5.2 Why Single-Patch Calibration Fails
@@
 This “anatomical dominance” arises from multiple physiological sources.
-The T-zone (forehead and glabella) exhibits higher sebaceous gland density than cheek regions, resulting in elevated sebum secretion that alters surface refractive index.[Man2009, Ohshima2013] Sebum forms a thin lipid film that increases specular reflection and modifies the balance between surface gloss and diffuse reflectance.[Baranoski2004] Under directional illumination, this regional variation produces angle-dependent color shifts.
-Facial skin thickness varies substantially: the periorbital region is among the thinnest (0.5–1.0 mm), while chin and forehead are thicker (1.5–2.0 mm).[Nouveau2016, Pioro2021] Since skin color arises from chromophore absorption at specific depths, thickness variations alter the optical path length and relative contributions of melanin versus hemoglobin.
+The T-zone (forehead and glabella) exhibits higher sebaceous gland density than cheek regions, resulting in elevated sebum secretion that alters surface optical properties.[Man2009] Sebum functions as an index-matching fluid ($n\approx 1.46$), reducing the refractive index mismatch at the stratum corneum ($n\approx 1.55$) via the optical clearing effect.[Tuchin2000, Ding2006] This reduction in refractive index discontinuity decreases surface scattering, allowing more photons to penetrate deeper into the epidermis and dermis where they interact with melanin and hemoglobin—producing “richer” or “darker” apparent skin color (lower $L^{*}$, higher chroma).
+Simultaneously, the smooth liquid sebum surface produces intense specular highlights under directional illumination.[Baranoski2004] These regional optical variations produce angle-dependent color shifts that a single CCM cannot correct.
+Facial skin thickness varies substantially: the periorbital region is among the thinnest (0.5–1.0 mm), while chin and forehead are thicker (1.5–2.0 mm).[Chopra2015] Since skin color arises from chromophore absorption at specific depths, thickness variations alter the optical path length and relative contributions of melanin versus hemoglobin.
 The chin region, exhibiting the largest color differences (mean $\Delta E=9.63$), presents a particularly challenging case due to pronounced surface curvature creating geometric variations in illumination.
@@
 Clinical applications must either: (a) avoid ratio-based indices, (b) develop biomarker-aware calibration that explicitly minimizes ITA error, or (c) employ uncertainty quantification to flag unreliable measurements.
-Third, the statistical significance of gender effects in our ANOVA ($p<0.001$) suggests that camera ISPs may exhibit differential behavior across demographics.
+Third, interestingly, gender also showed a statistically significant difference ($p<0.001$).
+Although its influence ($\eta^{2}=0.007$) is small compared to facial region, this suggests that camera ISPs may exhibit differential behavior across demographics.
 If “universal” CCMs are derived from specific populations, they may embed demographic biases—a fairness concern for global health applications.[Kinyanjui2020, Groh2021]
@@
 This study establishes that standard colorimetric calibration in smartphone dermatology creates false confidence: perceptually accurate images can yield unreliable clinical biomarkers.
-Using 43,425 images from 965 Korean subjects across DSLR, tablet, and smartphone devices, we demonstrate a fundamental “color-clinical decoupling” phenomenon.
+Using 43,424 images from 965 Korean subjects across DSLR, tablet, and smartphone devices, we demonstrate a fundamental “color-clinical decoupling” phenomenon.
 Our key findings are:
@@
 […]
+  3. 3.
+    Calibration errors disproportionately affect darker skin: ITA sensitivity to $b^{*}$ errors increases from $\sim$1.7°/unit for light skin (Type II/III) to >>4.5°/unit for dark skin (Type VI).
+A 2-unit calibration residual produces a 3.4° ITA shift in light skin but a 9.2° shift in dark skin—sufficient to cause skin type misclassification.
+This “metrological bias” means current calibration protocols fail to protect the most vulnerable populations from measurement error.
 Additionally, this work provides the first large-scale multi-device benchmark specifically for Fitzpatrick III–IV skin tones, addressing the “forgotten middle” in algorithmic fairness research.